A piece of wire 30 m long is cut into two pieces. One piece is bent into a square and the other is bent into a circle.


How much wire (in m) should be used for the square in order to maximize the total area?


M = 0


How much wire (in m) should be used for the square in order to minimize the total area?


M= ?


I've tried 16.8 but its wrong.


1
Expert's answer
2022-05-19T19:21:15-0400

A piece of wire, 30 m long, is cut in two sections: a and b. Then, the relation between a and b is:a+b=30

b=30-a

The section "a" is used to make a square and the section "b" is used to make a circle.


The section "a" will be the perimeter of the square, so the square side will be:


l=a/4


Then, the area of the square is:

"A_s=l^2=(a\/4)^2=a^2\/16"

The section "b" will be the perimeter of the circle. Then, the radius of the circle will be:

"2\pi r=b=30-a"

"r=\frac{30-a}{2\pi}"

The area of the circle will be:

"A_c=\pi r^2=\pi (\frac{30-a}{2\pi})^2=\frac{900-60a+a^2}{4\pi}"

The total area enclosed in this two figures is:

"A=A_s+A_c=\frac{a^2}{16}+\frac{900-60a+a^2}{4\pi}"

To calculate the extreme values of the total area, we derive and equal to 0:

"dA/da=\frac{2a}{16}+\frac{-60+2a}{4\pi}=\frac{a}{8}+\frac{a-30}{2\pi}=\frac{a\pi +4a-120}{8\pi}=0"

"a\pi+4a-120=0"

7.14a=120

a=16.8

We obtain one value for the extreme value, that is a=16.8.


We can derive again and calculate the value of the second derivative at a=16.8 in order to know if the extreme value is a minimum (the second derivative has a positive value) or is a maximum (the second derivative has a negative value):

"d^2A/da^2=\frac{\pi +4}{8\pi}>0"


As the second derivative is positive at a=16.8, this value is a minimum.


In order to find the maximum area, we analyze the function. It is a parabola, which decreases until a=16.8, and then increases.


Then, the maximum value has to be at a=0 or a=30, that are the extremes of the range of valid solutions.


When a=0 (and therefore, b=30), all the wire is used for the circle, so the total area is a circle, which surface is:

"A=\pi r^2=\pi (\frac{30}{2\pi})^2=71.62"

When a=30, all the wire is used for the square, so the total area is:

"A=a^2\/16=900\/16=56.25"


The maximum value happens for a=0.